Solve using the elimination method. -7x – y = -30 4x + y = 21

Answer by algebrahouse.com(1079)

-7x - y = -30
4x + y = 21
-------------
-3x = -9 {added the two equations together, eliminating the y}
x = 3 {divided both sides by -3}

4x + y = 21 {original bottom equation}
4(3) + y = 21 {substituted 3, in for x, into 4x + y = 21}
12 + y = 21 {multiplied 4 by 3}
y = 9 {subtracted 12 from both sides}

x = 3 and y = 9
For more help like this, ask the house: www.algebrahouse.com